OpenStudy (anonymous):

Solve y over negative 2− 3 = 12

OpenStudy (anonymous):

thats not an answer @uri

OpenStudy (uri):

I'm asking if that's the question.

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

-21?

OpenStudy (anonymous):

30 -30 18 -18 @uri those are the choices

OpenStudy (texaschic101):

(y/-2) - 3 = 12 ---- multiply everything by -2 to get rid of the fraction y + 6 = - 24 --- subtract 6 from both sides y = -24 - 6 y = -30

OpenStudy (anonymous):

thanks @texaschic101

OpenStudy (texaschic101):

any time you have fractions in equations or inequalities, you can get rid of them by multiplying everything in the problem by the common denominator.